OpenStudy (anonymous):

State whether the function is continuous at the indicated point. 36-9x/x-4 ; c=4

OpenStudy (anonymous):

36-9x/x-4=-9(x-4)/(x-4)=-9 so this function has a constant value for any value of x so it is continuous

OpenStudy (anonymous):

thank you!